# Flow-Induced Phase Separation for Active Brownian Particles in Four-Roll-Mill Flow

### 28 Oct 2025 v1

@@ III.3 Number Fluctuations
Figure 3(a) shows the giant number fluctuations of ABPs in the absence of any background flow for packing fractions ranging from $\rm\phi=0.1$ to $\rm 0.9$. At low densities ($\rm\phi<0.4$), we observe $\rm\alpha\approx 0.5$ in Fig. 3(c)(shown with red triangles), consistent with homogeneous distribution, which implies that there is no phase separation. As the system enters the MIPS regime at higher densities ($\rm\phi\geq 0.4$), the exponent increases towards $\rm\alpha\approx 1$ as shown in Fig. 3(c), indicating the emergence of large-scale density inhomogeneities and clustering driven by persistent self-propulsion [14].
In the presence of structured background flow, as shown in Fig. 3(b), for $\rm\phi\leq 0.4$, fluctuations remain normal with $\rm\alpha\approx 0.5$ clearly noticed in Fig. 3(c) with blue squares. However, beyond $\rm\phi>0.5$, the exponent increases toward unity, marking the onset of FIPS, where the combined effects of activity, crowding, and external flow drive the formation of dense, dynamic clusters [20].
As observed, MIPS emerges at relatively lower packing fractions $\rm\phi$ compared to the FIPS scenario, since the background flow enhances particle mobility and delays clustering. To quantify this difference, we plot the scaling exponent $\rm\alpha$ of the number fluctuations for both cases, without (red triangles) and with (blue squares) background flow. Figure 3(c) clearly illustrates the distinct behavior of $\rm\alpha$ as a function of $\rm\phi$ in the two situations, highlighting the effect of flow on the onset of clustering. Additionally, as shown in Fig. 3(c), the red triangles corresponding to the MIPS case attain higher values of the exponent $\rm\alpha$ compared to the blue squares representing the FIPS case. This indicates that number fluctuations are more pronounced in MIPS than in FIPS.
(a)Refer to caption(b)
Figure 6: (a) PDF of the Okubo–Weiss parameter $\rm\Lambda$ for different $\rm\phi\in[0.1,0.9]$ values. The emergence of asymmetry in the distribution with increasing $\phi$ indicates growing localization in strain-dominated regions. (b) Plot of skewness ($\rm\frac{\langle\Lambda^{3}\rangle}{\langle\Lambda^{2})\rangle^{3/2}}$) versus $\phi$.
III.4 Cluster Size Distribution
@@
We observed that the exponent $\beta$ for $\phi\geq 0.4$, as shown in Fig. 5(a), exhibits a clear difference between MIPS (red triangles) and FIPS (blue squares). The exponent $\beta$ is consistently higher in the presence of background flow, indicating the formation of smaller clusters due to the flow. In addition, we also observed the normalized maximum cluster size $\rm n_{max}/N$ across all values of $\phi$, where $\rm n_{max}$ is the largest cluster size (containing $\rm n$ particles), and $\rm N$ is the total number of particles corresponding to given packing fractions. As shown in Fig. 4(b), there is clear evidence that the presence of flow suppresses cluster formation.
III.5 PDF of Okubowise parameter
@@
Figure 6 presents the PDF of $\Lambda$ sampled along particle trajectories for varying particle packing fractions $\phi$. At low packing fractions ($\phi\lesssim 0.4$), the PDF is nearly symmetric about $\Lambda=0$, indicating a relatively uniform distribution of particles between vortical and strain-dominated regions. This suggests minimal preferential sampling by the particles at low concentrations.
However, as the packing fraction increases ($\phi\gtrsim 0.5$), the PDF becomes noticeably asymmetric, exhibiting a pronounced tail for negative values of $\Lambda$. This negative skewness indicates an increased likelihood of particles residing in strain-dominated regions of the flow, as shown in Fig. 6(b).
Figure 6(b) reveals an interesting trend: beyond $\phi=0.7$, the magnitude of skewness shows a slight decrease, which can be attributed to increased particle crowding. As the packing fraction rises, the finite size of particles forces a larger fraction of them into the vortical regions of the flow, as also evident from the snapshots in Appendix Fig. 7. This redistribution reduces the overall skewness magnitude. Nevertheless, the skewness remains negative across all densities, indicating that most particles continue to occupy the strain-dominated regions of the flow.
@@ IV Conclusion and Discussion
Our results reveal a clear crossover from a homogeneous phase at low densities to a distinct FIPS regime at higher densities ($\rm\phi\geq 0.6$). Unlike classical MIPS, which emerges due to self-trapping in crowded environments, FIPS arises from the interplay between particle motility, interparticle interactions, and external flow-induced vorticity. The emergence of persistent clusters in strain-dominated regions, stabilized by the background flow, characterizes this FIPS phase.
Through mean-square displacement (MSD) analysis, we observed a non-trivial crossover from ballistic to diffusive behavior with an intermediate plateau or bump, particularly in the FIPS regime, indicating transient trapping and clustering. The drift velocity of the ABPs remained nearly constant across $\rm\phi$, mainly driven by the steady flow field, while the effective diffusivity decayed quadratically with $\rm\phi$, consistent with enhanced clustering and reduced mobility at higher densities.
@@
Our findings open up several avenues for future exploration. Incorporating thermal fluctuations would help assess the robustness of FIPS under realistic conditions, while extending the system to three dimensions may reveal more intricate clustering behavior. Allowing feedback from ABPs to the flow field, i.e., considering hydrodynamic interactions, would move the model closer to experimental systems like bacterial suspensions. Exploring other flow topologies, such as shear or turbulent-like flows, could uncover new phases or transitions. Furthermore, studying anisotropic particles or alternative interaction potentials might broaden the observed collective behaviors. Finally, constructing effective theoretical models and comparing predictions with experimental realizations in microfluidic or colloidal systems could significantly advance our understanding of the flow-induced phase separation (FIPS).
V Appendix

### 14 Mar 2026 v2

@@ III.4 Number Fluctuations
Fig. 3(a) shows the giant number fluctuations of ABPs in the absence of any background flow for packing fractions ranging from $\rm\phi=0.1$ to $\rm 0.9$. At low densities ($\rm\phi<0.4$), we observe $\rm\alpha\approx 0.5$ (see the inset of Fig. 3(a)), consistent with homogeneous distribution, which implies that there is no phase separation. As the system enters the MIPS regime at higher densities ($\rm\phi\geq 0.4$), the exponent increases towards $\rm\alpha\approx 1$, indicating the emergence of large-scale density inhomogeneities and clustering driven by persistent self-propulsion [6].
In the presence of structured background flow, for $\rm\phi\leq 0.4$, the system exhibits equilibrium-like number fluctuations, $\rm\alpha\approx 0.5$ clearly indicated in the inset of Fig. 3(b). However, beyond $\rm\phi>0.5$, the exponent increases toward unity, marking the onset of FIPS, where the combined effects of activity, crowding, and external flow drive the formation of dense, dynamic clusters [20]. Here, The inset of Fig. 3(a) and 3(b) shows the scaling exponent as a function of $\phi$ for different system sizes ranging from $880$ to $10^{4}$. For sufficiently large particle numbers, giant number fluctuations are evident, providing reliable statistics for both MIPS and FIPS case.
Furthermore, MIPS emerges at relatively lower packing fractions $\rm\phi$ compared to the FIPS scenario. The background flow enhances particle mobility, which suppresses or delays clustering. As shown in Fig. 3(c), the MIPS case exhibits systematically larger values of $\rm\alpha$ than the FIPS case, indicating stronger number fluctuations and more pronounced density inhomogeneities in the absence of flow. The onset of giant number fluctuations is significantly sharper in the MIPS case, whereas in FIPS the increase of the fluctuation exponent occurs more gradually, as shown in Fig. 3(c). This qualitative difference suggests a discontinuous-like transition in MIPS and a more continuous-like transition in FIPS. However, this inference is based solely on the behavior of the number-fluctuation exponent; a rigorous characterization of the transition and its order lies beyond the scope of the present work and will be addressed in a future study.
(a)(b)(c)(d)
Figure 4: Normalized cluster size distribution $\rm P(n)/P(1)$ as a function of ratio of cluster size to maximum number of particle at corresponding packing fractions $\rm n/N_{max}$ for ABPs with size $\rm a=0.02$ at various packing fractions $\rm\phi\in[0.1,0.9]$: (a) without background flow exhibiting MIPS behavior, and (b) with background flow exhibiting FIPS behavior. (c) Exponent of the cluster size distribution (CSD) and (d) normalized maximum cluster size shown as functions of packing fraction $\phi$, comparing systems in the absence and presence of background flow.
III.5 Cluster Size Distribution
@@
We observed that the exponent $\beta$ for $\phi\geq 0.4$, as shown in Fig. 4(c), exhibits a clear difference between MIPS (red asterisk) and FIPS (blue squares). The exponent $\beta$ is consistently higher in the presence of background flow, indicating the formation of smaller clusters due to the flow. In addition, we also observed the normalized maximum cluster size $\rm n_{max}/N$ across all values of $\phi$, where $\rm n_{max}$ is the largest cluster size (containing $\rm n$ particles), and $\rm N$ is the total number of particles corresponding to given packing fractions. As shown in Fig. 4(b), there is clear evidence that the presence of flow suppresses cluster formation.
III.6 PDF of Okubo-Weiss parameter
@@
Fig. 5 presents the PDF of $\Lambda$ sampled along particle trajectories for varying particle packing fractions $\phi$. At low packing fractions ($\phi\lesssim 0.4$), the PDF is nearly symmetric about $\Lambda=0$, indicating a relatively uniform distribution of particles between vortical and strain-dominated regions. This suggests minimal preferential sampling by the particles at low packing fraction. However, as the packing fraction increases ($\phi\gtrsim 0.5$), the PDF becomes noticeably asymmetric, exhibiting a pronounced tail for negative values of $\Lambda$. This negative skewness indicates an increased likelihood of particles residing in strain-dominated regions of the flow, as shown in inset of Fig. 5.
The inset of Fig. 5 reveals an interesting trend: beyond $\phi=0.7$, the magnitude of skewness shows a slight decrease, which can be attributed to increased particle crowding. As the packing fraction rises, the finite size of particles forces a larger fraction of them into the vortical regions of the flow, as also evident from the snapshots in Appendix Fig. 6. This redistribution reduces the overall skewness magnitude. Nevertheless, the skewness remains negative across all densities, indicating that most particles continue to occupy the strain-dominated regions of the flow.
Figure 5: PDF of the Okubo–Weiss parameter $\rm\Lambda$ for different $\rm\phi\in[0.1,0.9]$ values. The emergence of asymmetry in the distribution with increasing $\phi$ indicates growing localization in strain-dominated regions. Inset: Plot of skewness ($\rm\frac{\langle\Lambda^{3}\rangle}{\langle\Lambda^{2})\rangle^{3/2}}$) versus $\phi$.
@@ IV Conclusion and Discussion
Our results reveal a clear crossover from a homogeneous phase at low densities to a distinct FIPS regime at higher densities ($\rm\phi\gtrsim 0.5$). Unlike classical MIPS, which emerges due to self-trapping in crowded environments, FIPS arises from the interplay between particle motility, interparticle interactions, and external flow-induced vorticity. The emergence of persistent clusters in strain-dominated regions, stabilized by the background flow, characterizes this FIPS phase.
Through mean-square displacement (MSD) analysis, we observed a non-trivial crossover from ballistic to diffusive behavior with an intermediate plateau, particularly in the FIPS regime, indicating transient trapping. The drift velocity of the ABPs remained nearly constant across $\rm\phi$, mainly driven by the steady flow field, while the effective diffusivity decayed quadratically with $\rm\phi$, consistent with enhanced clustering and reduced mobility at higher densities.
The overlap function provides further insight into the origin of the plateau observed in the mean squared displacement. It indicates that periodic flow induces transient caging, which is temporarily disrupted by activity and followed by transient recaging due to crowding in strain-dominated regions. The relaxation time exhibits an exponential decay in both cases; however, its dependence on packing fraction differs between the MIPS and FIPS regimes. While particle interactions primarily drive the dynamical slowdown, the imposed flow redistributes particles in the system, leading to a more gradual development of caging and clustering.
@@
Our findings open up several avenues for future exploration. Incorporating thermal fluctuations would help assess the robustness of FIPS under realistic conditions, while extending the system to three dimensions may reveal more intricate clustering behavior. Allowing feedback from ABPs to the flow field, i.e., accounting for hydrodynamic interactions, would move the model closer to experimental systems such as bacterial suspensions. Exploring other flow topologies, such as shear or turbulent-like flows, could uncover new phases or transitions. Furthermore, studying anisotropic particles or alternative interaction potentials might broaden the observed collective behaviors. Finally, constructing effective theoretical models and comparing predictions with experimental realizations in microfluidic or colloidal systems could significantly advance our understanding of the flow-induced phase separation (FIPS).
V Acknowledgment
SDP acknowledges Mustansir Barma for a fruitful discussion and suggestions. SDP acknowledges the receipt of University Grants Commission (UGC) 1608(CSIRNETJUNE2019), which supported this study. AG acknowledges the funding from SERB-India (grant no. MTR/2022/000232, grant no. CRG/2023/007056-G); DST-India (grant no.DST/NSM/R&D HPC Applications/2021/05 and grant no. SR/FST/PSI-309 215/2016). KS acknowledges the DST-INSPIRE Grant (No: DST/INSPIRE Fellowship/[IF230178]) for financial support.
VI Appendix